BALANCE SHEETS: Assets:                                       &n

BALANCE SHEETS:

Assets:

                                                                                   2017                            2018

                        Cash                                                  74,181                         66,301

                        Accounts Receivable                           35,673                         48,995

                        Inventory                                            4,855                           3,986

                        Other Current Assets                          13,936                         12,057

                        Fixed Assets, net                                 33,783                         41,304

                        Investments                                         212,891                       233,082

                        Total Assets                                        375,319                       405,725

Liabilities and Equity:

                        Accounts Payable                             44,242                         55,888

                        Other Current Liabilities                     50,226                         55,416

                        Long-Term Debt                                 103,703                       102,519

                        Other Noncurrent Liabilities             43,251                         48,209

                        Common Stock                                   35,567                         33,293

                        Retained Earnings                               98,330                         110,400

                        Total Liabilities and Equity                  375,319                       405,725

INCOME STATEMENT:

                                                                                                            FY 2018

            Revenue                                                                                  265,595

            Cost of Goods Sold                                                                163,756

            General and Administrative                                                       14,793

            Depreciation Expense                                                             10,903

            Earnings Before Interest and Taxes                                         76,143

            Interest Expense                                                                      3,240

            Pretax Net Income                                                                  72,903

            Income Taxes                                                                          13,372

            Net Income                                                                             59,531

  1. What was Apple’s dividend payout ratio during fiscal 2018?
Answer #1

Dividends paid in fiscal 2018=98330+59531-110400=47461
Dividend payout ratio=47461/59531=79.72%